13) Find the coordinates of the other endpoint given a midpoint (½, 5) and endpoint (-5, -1).
Answer by stanbon(75887)

Find the coordinates of the other endpoint given a midpoint (½, 5) and endpoint (-5, -1).
-----
Let the other end point be (x,y). Solve for "x" and for "y":
------
(-5 + x)/2 = 1/2
-5+x = 1
x = 6
----
(-1+y)/2 = 5
-1+y = 10
y = 11
-------
Other end point is (6,11)
================================
Cheers,
Stan H.